OpenStudy (anonymous):

find the slope of a line parallel to each given line?? y= 1/4x + 2

OpenStudy (anonymous):

equation of a linear graph: y=mx+c, where: m=gradient of the line and c= intersection point with y-axis since u have y=(1/4)*x+2, therefore m=1/4.

OpenStudy (anonymous):

o okay thank u so any parallel line would be 1/4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

parallel lines have the same slope so SINCE the slope of the one you were given is 1/4 then the slope of any line parallel to it would be 1/4.

OpenStudy (anonymous):

You statement "so any parallel line would be 1/4" - only in this problem. Just making you are sure on this.

OpenStudy (anonymous):

o okay thank you

OpenStudy (anonymous):

do you know what the slope of a line perpendicular to x= -2 would be

OpenStudy (anonymous):

if the line hav slope equal to m, the perpendicular line have slope -1/m